OpenStudy (anonymous):

How to assign classes in Grouped Frequency Distribution in Elementary Statistics?

OpenStudy (campbell_st):

each class needs to be the same width... e.g 0-9, 10.19.20-29 etc... within each class you need a class centre... which is the end poinst for each class/2 e.g. (0+9)/2 = 4.5, (10 + 19)/2 = 14.5 etc